OpenStudy (anonymous):

-2 <= 3-5x <= 18 . I need help solving this question . Any help is much appreciated .

geerky42 (geerky42):

You would need to isolate x. What to do?

geerky42 (geerky42):

What to do about 3 in 3-5x?

OpenStudy (koikkara):

Carnifex...hey Before you go Deep, Welcome to Open Study !!

OpenStudy (texaschic101):

-2 < = 3 - 5x <= 18 -- subtract 3 from all sides -2 - 3 <= 3 - 3 - 5x <= 18 - 3 -- combine like terms -5 <= -5x <= 15 -- now divide by -5, and change the inequality signs -5/-5 <= -5x/-5 <= 15/-5 1 >= x >= -3 any questions ?

OpenStudy (texaschic101):

or you could have also done it by seperating them.. -2 <= 3 - 5x -2 - 3 <= -5x -5 <= -5x 1 >= x 3 - 5x <= 18 -5x <= 18 - 3 -5x <= 15 x >= -3 leaving us with : -3 <= x <= 1

OpenStudy (texaschic101):

either way is correct

OpenStudy (anonymous):

thank you so much